OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = –3/4,x – 2?

OpenStudy (anonymous):

the given graph is a line.... slope of the given line is -3/4 slope of a line perpendicular to a given line is ... \[-\frac{1}{-\frac{3}}{4}\] which is equal to 4/3 using point slope for to find equation of a line....given point is (12,4) y-4=(4/3)(x-12) 3y-12=4x-48 4x-3y=36 is the resulting line